find x so that x+2,4x and 6x+1 are terms of an arithmetic sequence
Answer by dkppathak(439)

find x so that x+2,4x and 6x+1 are terms of an arithmetic sequence
solution

if a ,b ,c are in AP than we know that b=a+c/2 or 2b=a+c
by same application
2*4x=x+2+6x+1
8x=7x+3
8x-7x=3
x=3